Operations Report – December 31st

Crew 218 Operations Report 31-Dec-19
SOL: 09

Name of person filing report: Luz Maria Agudelo Urrego

Non-nominal systems: NA

Generator: run
Hours run: 14
From what time last night: 1730
To what time this morning: 0730
List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before the generator is run at night: 71%

Diesel Reading – 75%
Station Propane Reading – 66%
Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 50
Water Meter (units): 0147175.1
Water (static tank) (gal): 300
Static to Loft Pump used – No
Water in Green Hab (gal): 88
Water in Science Dome (gal): 0

Toilet tank emptied: No

Deimos rover used: Still in the workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: Assigned to director
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: EVA 6
Hours: 120.6
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 63%
Currently charging: Yes

Spirit rover used: EVA 7
Hours: 120.9
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 87%
Currently charging: Yes
Opportunity rover used: Still in the workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:

Curiosity rover used: EVA 6
Hours: 125.9
Beginning charge: 100%
Ending charge: (On return from EVA, before charging): 71%
Currently charging: Yes
Curiosity rover used: EVA 7
Hours: 126.2
Beginning charge: 100%
Ending charge: (On return from EVA, before charging): 87%
Currently charging: Yes

Notes on rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use: N/A
Oil Added? No
# Hours the ATVs were used today: 0
Notes on ATVs: N/A

HabCar used, and why, where? No
CrewCar used, and why, where? No

General notes and comments: N/A

Summary of the internet: Nominal
Summary of suits and radios: Nominal
Summary of Hab operations: The pipe from the static tank through the loft tank is still frozen. The crew filled out the Loft Tank by carrying water buckets from the Static Tank again.
Summary of Science Dome operations: Nominal
Summary of RAM operations: Nominal
Summary of any observatory issues: Nominal
Summary of health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: NA

Journalist Report – December 31st

Journalist Report
Ben Durkee, Crew 218 Journalist

Sol 09

Today began like any other day on Mars: awoken dim and early by the blasting of a select song from the Futurama soundtrack. We stirred ourselves some hot chocolate – the only dehydrated food we seem to have an abundance of – and gathered around the window to watch the sunrise. Once the sun peeked out from its azimuthal hiding spot, it seemed we were already in a rush to eat breakfast and prepare for our first EVA of the day.

With this being our 6th EVA, we were already pros at the process. Myself, Shefali, and Cesare were suited up, depressurized, and revving up the rovers in 15 minutes flat. Our goal for this EVA was to head south of the Hab. Not for warmer pastures, but for a signal analysis experiment. My research involves measuring ambient radio signals around the Hab, so we were testing whether or not the ridge between the Hab and our destination would have a dampening effect on such signals.

The ridge in question is called Kissing Camel Ridge, supposedly named because the rock formations look like two titular camels face to face. However, I adamantly believe that anyone who claims to see two camels there is either a fraud or suffering from terminal "Emperor’s New Clothes" Syndrome.

We accomplished our own personal research goals in the area and spent our remaining time capturing pictures of the grandiose landscape. Admittedly also a few (hundred) pictures of each other in fabulous poses. We loaded back up the rovers with our cumbersome equipment and slogged back to base. The roads today are swampy concoctions of mud and slushy ice under the facade of neatly laid snow. We returned home cold, muddy, and accomplished.

As soon as we escaped our life-saving mobile prisons (that we are very grateful for), we tagged in the next EVA crew. Pat, seismologist extraordinaire, and LuzMa, dehydrated mango connoisseur performed an engineering maintenance EVA while they waited for the rovers to recharge. Once T and Custy had drunk their fill of our precious electricity, they valiantly carried the two Marstronauts to their goal. That goal being: south. Just slightly farther south than we had ventured. Always one-upping us, those two.

After performing a long series of geophonic tests that we Hab-dwellers could intermittently hear over the radios, they too braved the earthy sludge on their homeward bound. They got back to breathable air just in time to join the sequel to yesterday’s fire brigade! We had waited a majority of the day to ferry water in the hopes that the pipes would thaw, but no dice. On the bright side, since we had used some of yesterday’s water to wash the dishes, we had a multitude of clean pots and other water receptacles for our thirst-fueled operation.

1000 rotations of dumping pots and slamming my head into the low ceiling later, our water was at a more than acceptable level. We filled it up a bit extra so that people had the opportunity to shower before our New Year’s festivities if desired. A few people seized the hygienic opportunity, and then we spontaneously dove into a viewing of Shrek. As a neutral journalistic observer, I must report that tensions are beginning to rise. There may be a formation of factions between those who want to watch Mulan next and those who would prefer Hot Fuzz. The first Martian Civil War breweth.

The other thing beginning to rise is our rosemary bread for the great New Year’s feast! Very soon we will put aside our differences and eat the vast majority of our remaining food supply. It’s strange to celebrate the anniversary of the Earth’s revolution around the sun from its planetary next-door neighbor. There’s no way I’d rather spend it than breaking Hab-made bread with my fellow crewmates.

What a fantastic way to usher in the new decade. Happy New Year, everyone!

Science Report – December 31st

Crew 218 Science Report 31-DEC-19
Crew Science Officer: Dr. Jonathan R. Buzan

1. Decision Making in support of autonomy for crew EVAs:

Hypothetical scenario involving a broken leg during EVA. EVA crew went through the emergency procedures.

2. Mars surface weather:

Measurements were collected during today’s EVA.

3. Subsurface structure on Mars:

Data from yesterday was sent to Purdue University for analysis. Seismic reflection mapping instrument was deployed in today’s EVA.

4. Detecting radio signal strength:

Field measurements were collected during today’s EVA. Yesterday’s data is undergoing analysis.

5. EVA workload analysis:

Survey data was collected for today’s EVAs.

6. EMU (Extravehicular Mobility Unit) ergonomic assessment:

Survey data was collected for today’s EVAs.

7. Environmental Stresses over MDRS habitat and Crew Members and projection over Martian Terrain:

Measurements were expanded from stateroom to habitat floor.

8. Messier and other space objects for outreach:

Cloudy: nothing to report.

9. Reliability and maintenance:

No EVA: nothing to report.

10. Medical readings in preparation for the future crew-wide project:

The Zephyr bio-monitor was worn all day yesterday.

11. Collection of clay, shale, and hematite samples:

6 rock samples were collected.

12. Media and outreach:

Nothing to report.

Glassware check out: None

A reminder to all crewmembers: There’s a $300 fine for using any glassware material without MDRS Mission support permission.

EVA Report – December 31st

EVA #: 6

Author: Cesare Guariniello

Purpose of EVA: weather observation project; EVA and EMU evaluation project

Start time: 9:57
End time: 12:01

Narrative: Cesare, Ben, and Shefali drove to the north side of Kissing Camel range, where Ben took radio measurements and Cesare proposed potential scenarios for crew decision-making.

After collecting a few rock samples, the crew drove the rovers to the south of the range, which is exposed and clear of snow.

More radio measurements were taken, and the crew could climb the south side of the West Kissing Camel range. On the way back, Ben took more radio measurements along the first ridge south of the habitat.

Destination: Kissing Camel

Coordinates (use UTM NAD27 CONUS): 12S 518300, 4249000

EVA Participants: Cesare Guariniello, CMD; Ben Durkee, JOU Shefali Rana, HSO

Road(s) and routes per MDRS Map: Cow Dung Road to the south of Kissing Camel Range

Mode of travel: Driving, then work in the field and walking

Vehicles Used: Curiosity and Spirit

EVA #: 7

Author: Pat Pesa

Purpose of EVA: crew decision-making evaluation and radio measurements (during transfers); seismic measurements to evaluate subsurface layers (once at destination); EVA and EMU evaluation (once back in habitat)

Start time: 12:49

End time: 14:57

Narrative: Luz Ma and Pat took the Curiosity and Spirit rovers to just south of Kissing Camel range to the Hippo formation. Pat set up his Seismic equipment with the help of Luz Ma (lots of untangling wires) and started recording data.
It is a quite fun experiment because, after the tedious setup, you get to "smack" the ground with a sledgehammer to generate waves for the geophones to pick up! After these tests, Luz Ma continued her Climate and Weather readings, and we embarked on a short journey near the area.

Besides the overabundance of slushy snow and mud, the EVA was greatly successful and led to some great science and exploration.

Destination: South of Kissing Camel along Cow Dung Rd

Coordinates (use UTM NAD27 CONUS): 12S 51900, 4248600

EVA Participants: Pat Pesa, GEO; Luz Ma Agudelo, ENG

Road(s) and routes per MDRS Map: Cow Dung Road to the northwest of Zubrin’s head

Mode of travel: Driving and walking

Vehicles Used: Curiosity and Spirit

Greenhab Report – December 31st

Crew 218 GreenHab Report 31-DEC-19

Crew GreenHab Officer: Dr. Jonathan R. Buzan and HSO Shefali Rana

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 23.4°C; 18%

10:00M

Floor Unit: 15°C
Electronic: 21.8°C
humidity 19%

3:00PM

Floor Unit: 21°C
Electronic: 28.5°C
humidity 16%

7:00PM

Floor Unit: 15°C
Electronic: 19.6°C
humidity 20%

Max: 29.5; 22%
Min: 19.6°C; 20%

Hours of supplemental light: Light system 6:00 PM-11:00 AM.

Daily water usage for crops: 4.9Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~88 Gal.

Time(s) of watering for crops:

10:00 AM. 1Gal. Watered snow peas and radishes.
7:00 PM. 3.7Gal.

SEASONAL PLANTING HISTORY:

Change to crops: None.

OTHER NOTES:

1st sprouts: N/A

—Mystery plant has a spinach vegetal taste.
—Hanging snow peas and icicle radishes are wilted. Used 1Gal. of water at 10 AM.

Harvest:

Basil: 6g
Rosemary: 4g

Commander Report – December 31st

Crew 218 Commander Report – 31-12-2019

Sol 9 – End of Earth year 2019

While we are preparing our feast to celebrate the last day of Earth year 2019, we are reviewing what was a good last day.

Finally, the Sun has been shining all day long and even if the snow did not completely melt, at least the EVAs in sub-freezing temperatures were much more pleasant than the past ones. In the afternoon, we started popping a few New Year wishes from Central Europe.

Due to pipes still being frozen, we had another chain of water upstairs, so as to be able to take some nice showers!

Now waiting for some guests from the nearby habitat to celebrate together.

Goodnight and Happy New Year!!

Cesare Guariniello, Commander Crew 218

Sol Summary – December 31st

Sol: 09

Summary Title: New Years New Mars

Author’s name: Pat Pesa

Mission Status: Crew is feeling festive!

Sol Activity Summary: Another beautiful Martian winter day that led to two very productive but muddy EVAs.

Look Ahead Plan: More EVAs! We only have a few days left and the weather looks great so we want to make the most of it.

Anomalies in work: None

Weather: Sunny, around 29 degrees

Crew physical status: Healthy

EVA: EVA #6 this morning went slightly north of kissing camel ridge for Radio Signal measurements. EVA #7 this afternoon went slightly south of kissing camel ridge to the formation formerly known as "Hippo" for the Seismic Survey for the Ground Structure Project and taking Climate data.

Reports to be filed: sol summary, commander report, operations report, GreenHab report, journalist report, EVA report, EVA request, science report

Support Requested: none

Pat Pesa
Geologist, MDRS 218